# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ B.1.1 Proof of Theorem 1
$\frac{10\gamma\rho}{\mathfrak{s}}=\log\left(1+\frac{p}{\mathfrak{s}^{2}}\log\left(\frac{3}{2\sqrt{2}}\right)+\sqrt{\left(1+\frac{p}{\mathfrak{s}^{2}}\log\left(\frac{3}{2\sqrt{2}}\right)\right)^{2}-1}\right),$ (B.2)
which together with $\mathfrak{s}\log\bigl{(}1+\tfrac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\bigr{)}\leq\sqrt{\tfrac{5}{2}\log\left(\tfrac{3}{2\sqrt{2}}\right)\tfrac{\Delta}{\tau}}$ implies
$2\gamma\rho^{2}\leq\frac{\mathfrak{s}\log\left(1+\frac{p}{\mathfrak{s}^{2}}\vee\frac{\sqrt{p}}{\mathfrak{s}}\right)\rho}{5}=\frac{\tau\mathfrak{s}^{2}\log^{2}\left(1+\frac{p}{\mathfrak{s}^{2}}\vee\frac{\sqrt{p}}{\mathfrak{s}}\right)}{5\Delta}\leq\frac{1}{2}\log\left(\frac{3}{2\sqrt{2}}\right).$
As in Part (i), we obtain $\alpha\leq 1/2$, and thus $\inf_{\widehat{\theta}}\;\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\;\mathsf{P}\left(\bigl{|}\widehat{\theta}-\theta(\mathsf{P})\bigr{|}\geq\gamma\right)\geq 1/4$. Also we can derive from (B.2) that
@@
$\frac{\gamma\rho}{\mathfrak{s}}\geq\frac{1}{5}\log\left(\frac{3}{2\sqrt{2}}\right)\log\left(1+\frac{p}{\mathfrak{s}^{2}}\vee\frac{\sqrt{p}}{\mathfrak{s}}\right).$
Note also that (B.2) implies $\gamma\leq\Delta/(10\tau)\leq\Delta/2$. ∎
@@ B.1.2 Supporting lemmas
For $\bm{u}\in\mathbb{R}^{p}$, let $\mathsf{P}_{\bm{u}}$ denote the centred Gaussian distribution with covariance matrix
$\begin{bmatrix}\mathbf{I}&\bm{u}\\ \bm{u}^{\top}&\bm{u}^{\top}\bm{u}+\sigma^{2}\end{bmatrix}\in\mathbb{R}^{(p+1)\times(p+1)}.$
If $|\bm{u}|_{2}=|\bm{v}|_{2}=\kappa<\sigma/\sqrt{2}$, then
$\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\frac{\mathrm{d}\mathsf{P}_{\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\frac{\mathrm{d}\mathsf{P}_{\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)=\left(\left(1-\frac{\langle\bm{u},\bm{v}\rangle}{\sigma^{2}}\right)^{2}-\left(\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}\right)^{-1/2}.$
@@
It follows from a general result for centred multivariate Gaussians (see e.g. Lemma 11 in Cai and Guo,, 2017). For completeness, we give below a direct computation:
@@
$=\mathrm{det}\left(\mathbf{I}-\frac{1}{\sigma^{2}}\left(\bm{u}\bm{v}^{\top}+\bm{v}\bm{u}^{\top}\right)\right)^{-1/2}$
$=\left(\left(1-\frac{\langle\bm{u},\bm{v}\rangle}{\sigma^{2}}\right)^{2}-\left(\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}\right)^{-1/2},$
where the last equality can be seen from the fact that
$\left(\mathbf{I}-\frac{1}{\sigma^{2}}\left(\bm{u}\bm{v}^{\top}+\bm{v}\bm{u}^{\top}\right)\right)(\bm{u}\pm\bm{v})=\left(1-\frac{\langle\bm{u},\bm{v}\rangle}{\sigma^{2}}\mp\frac{\kappa^{2}}{\sigma^{2}}\right)(\bm{u}\pm\bm{v}).$
@@
Note that in Lemma B.2 the condition $\kappa<\sigma/\sqrt{2}$ is only needed to ensure the invertibility of the matrix $\mathbf{I}-\frac{1}{\sigma^{2}}\left(\bm{u}\bm{v}^{\top}+\bm{v}\bm{u}^{\top}\right)$. In the special case of $\kappa^{2}+\langle\bm{u},\bm{v}\rangle=\sigma^{2}$, by the same calculation, we have
$\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\frac{\mathrm{d}\mathsf{P}_{\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\frac{\mathrm{d}\mathsf{P}_{\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)=\left(1-\frac{\langle\bm{u},\bm{v}\rangle}{\sigma^{2}}+\frac{\kappa^{2}}{\sigma^{2}}\right)^{-1/2}.$
However, if $\kappa^{2}+\langle\bm{u},\bm{v}\rangle>\sigma^{2}$, we have $\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\tfrac{\mathrm{d}\mathsf{P}_{\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\tfrac{\mathrm{d}\mathsf{P}_{\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)=\infty$.
@@
See either the first proof in Hoeffding, (1963, Theorem 4 on pages 21–23), or a more explicit one in Aldous, (1985, Proposition (20.6) on pages 171–172), or a systematic treatment in Kemperman, (1973, Section 3). ∎

### v3

@@ B.1.1 Proof of Theorem 1
$\mathsf{E}\left[\left(\left(1-\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\frac{\kappa^{2}}{\sigma^{2}2^{(m+\ell)/2}}\right)^{2}-\left(\frac{\kappa^{2}}{\sigma^{2}2^{(m+\ell)/2}}\right)^{2}\right)^{-2^{(m\wedge\ell)-1}}\right]$
$\leq\,$ $\mathsf{E}\left[\left(1-\frac{2\rho}{2^{(m+\ell)/2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}-\frac{\rho^{2}}{2^{m+\ell}}\right)^{-2^{(m\wedge\ell)-1}}\right]$
$\leq\,$ $\exp\left(\frac{\rho^{2}}{2^{m\vee\ell}}\right)\mathsf{E}\left[\left(1-\frac{2^{1-(m+\ell)/2}\rho}{1-2^{-(m+\ell)}\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)^{-2^{(m\wedge\ell)-1}}\right].$
By the symmetry of Rademacher distribution, we can find an event $A$ with $\mathsf{P}(A)=1/2$ such that $A\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\geq 0\}$ and $A^{c}\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\leq 0\}$. As $2^{1-(m+\ell)/2}\rho/(1-2^{-(m+\ell)}\rho^{2})\leq 3/4$, we apply again Lemma B.1 and obtain, for $m,\ell\in\mathcal{M}$,
$\mathsf{E}\left[\left(1-\frac{2^{1-(m+\ell)/2}\rho}{1-2^{-(m+\ell)}\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)^{-2^{(m\wedge\ell)-1}}\right]$
@@
$\leq\;$ $\mathsf{E}\left[\mathbb{I}_{A}\cdot\exp\left(2^{m\wedge\ell}\cdot\frac{2^{1-(m+\ell)/2}\rho}{1-2^{-(m+\ell)}\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)\right]$
$\qquad\quad+\mathsf{E}\left[\mathbb{I}_{A^{c}}\cdot\exp\left(\frac{2^{m\wedge\ell}}{4}\cdot\frac{2^{1-(m+\ell)/2}\rho}{1-2^{-(m+\ell)}\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)\right]$
$\leq\;$ $\left(\frac{1}{2}\mathsf{E}\left[\exp\left(2^{(m\wedge\ell)+1}\cdot\frac{2^{1-(m+\ell)/2}\rho}{1-2^{-(m+\ell)}\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)\right]\right)^{1/2}$
$\qquad\quad+\left(\frac{1}{2}\mathsf{E}\left[\exp\left(2^{(m\wedge\ell)-1}\cdot\frac{2^{1-(m+\ell)/2}\rho}{1-2^{-(m+\ell)}\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)\right]\right)^{1/2}$
$\leq\;$ $\sqrt{2}\left[\cosh\left(\frac{9\rho}{2^{|m-\ell|/2+1}\mathfrak{s}}\right)\right]^{|S\cap[\mathfrak{s}]|/2},$
where the second inequality is by Hölder’s inequality. Note that the summation involving $|S\cap[\mathfrak{s}]|$ can be seen as the expectation of the hypergeometric distribution with parameters $(p,\mathfrak{s},\mathfrak{s})$. Let $\xi\sim\mathrm{Bin}(\mathfrak{s},\mathfrak{s}/p)$. Applying Lemma B.3, we obtain
$\frac{1}{{p\choose\mathfrak{s}}}\sum_{S\subseteq[p],\,|S|=\mathfrak{s}}\left[\cosh\left(\frac{9\rho}{2^{|m-\ell|/2+1}\mathfrak{s}}\right)\right]^{|S\cap[\mathfrak{s}]|/2}$ $\leq\mathsf{E}\left[\cosh\left(\frac{9\rho}{2^{|m-\ell|/2+1}\mathfrak{s}}\right)^{\xi/2}\right]$
$=\left\{1+\frac{\mathfrak{s}}{p}\left(\cosh\left(\frac{9\rho}{2^{|m-\ell|/2+1}\mathfrak{s}}\right)^{1/2}-1\right)\right\}^{\mathfrak{s}}.$
Let $\bar{m}$ and $\bar{\ell}$ be independent uniform random variables on $\mathcal{M}$. Then,
$\chi^{2}(\mathsf{P}^{(1)},\mathsf{P}^{(0)})+1\leq\sqrt{2}\mathsf{E}\left[\exp\left(\frac{\rho^{2}}{2^{\bar{m}\vee\bar{\ell}}}\right)\left\{1+\frac{\mathfrak{s}}{p}\left(\cosh\left(\frac{9\rho}{2^{|\bar{m}-\bar{\ell}|/2+1}\mathfrak{s}}\right)^{1/2}-1\right)\right\}^{\mathfrak{s}}\right].$
Note that, for $m,\ell\in\mathcal{M}$, $\sqrt{2}\exp\bigl(\rho^{2}/2^{m\vee\ell}\bigr)\leq\sqrt{2}\exp\bigl(2\underline{c}_{1}^{2}\bigr)\leq 3/2.$ Next we consider separately two scenarios:
Dense regime:
$s\geq\sqrt{p\log\log(n)}$. It implies that $\rho\leq c_{1}\sqrt{p\log\log(n)}$. By the basic inequalities $\cosh(x)\leq\exp(x^{2}/2)$, $\exp(x)-1\leq x\exp(x)$, and $1+x\leq\exp(x)$ for $x\geq 0$, we have
$\mathsf{E}\left[\left\{1+\frac{\mathfrak{s}}{p}\left(\cosh\left(\frac{9\rho}{2^{|\bar{m}-\bar{\ell}|/2+1}\mathfrak{s}}\right)^{1/2}-1\right)\right\}^{\mathfrak{s}}\right]\leq\mathsf{E}\left[\exp\left\{\frac{81\rho^{2}}{p2^{|\bar{m}-\bar{\ell}|+4}}\exp\left(\frac{81\rho^{2}}{2^{|\bar{m}-\bar{\ell}|+4}\mathfrak{s}^{2}}\right)\right\}\right].$
We split the expectation above into two parts corresponding to $|\bar{m}-\bar{\ell}|\leq\log\log(n)$ and $|\bar{m}-\bar{\ell}|>\log\log(n)$. Then, the first part satisfies
$\mathsf{E}\left[\exp\left\{\frac{81\rho^{2}}{p2^{|\bar{m}-\bar{\ell}|+4}}\exp\left(\frac{81\rho^{2}}{2^{|\bar{m}-\bar{\ell}|+4}\mathfrak{s}^{2}}\right)\right\}\mathbb{I}_{\left\{|\bar{m}-\bar{\ell}|\leq\log\log(n)\right\}}\right]$
$\leq\;$ $\exp\left\{\frac{81\rho^{2}}{16p}\exp\left(\frac{81\rho^{2}}{16\mathfrak{s}^{2}}\right)\right\}\mathsf{P}\left\{|\bar{m}-\bar{\ell}|\leq\log\log(n)\right\}$
$\leq\;$ $\log(n)^{\frac{81\underline{c}_{1}^{2}}{16}\exp\left(\frac{81\underline{c}_{1}^{2}}{16}\right)}\cdot\frac{2\log\log(n)+1}{(1-2\nu)\log_{2}(n)-2}$
which goes to zero as $n\to\infty$, as $\frac{81\underline{c}_{1}^{2}}{16}\exp\left(\frac{81\underline{c}_{1}^{2}}{16}\right)<\frac{2}{5}$. For the second part, we have
$\mathsf{E}\left[\exp\left\{\frac{81\rho^{2}}{p2^{|\bar{m}-\bar{\ell}|+4}}\exp\left(\frac{81\rho^{2}}{2^{|\bar{m}-\bar{\ell}|+4}\mathfrak{s}^{2}}\right)\right\}\mathbb{I}_{\left\{|\bar{m}-\bar{\ell}|>\log\log(n)\right\}}\right]$
$\leq\;$ $\exp\left\{\frac{81\rho^{2}}{p2^{\log\log(n)+4}}\exp\left(\frac{81\rho^{2}}{16\mathfrak{s}^{2}}\right)\right\}\mathsf{P}\left\{|\bar{m}-\bar{\ell}|>\log\log(n)\right\}$
$\leq\;$ $\exp\left\{\frac{81\underline{c}_{1}^{2}\log\log(n)}{16\log(n)^{\log(2)}}\exp\left(\frac{81\underline{c}_{1}^{2}}{16}\right)\right\},$
which goes to one, as $n\to\infty$.
Sparse regime:
$s<\sqrt{p\log\log(n)}$. By the basic inequalities $\cosh(x)\leq\exp(x)$ and $1+x\leq\exp(x)$ for $x\geq 0$, we have
$\mathsf{E}\left[\left\{1+\frac{\mathfrak{s}}{p}\left(\cosh\left(\frac{9\rho}{2^{|\bar{m}-\bar{\ell}|/2+1}\mathfrak{s}}\right)^{1/2}-1\right)\right\}^{\mathfrak{s}}\right]\leq\mathsf{E}\left[\exp\left(\frac{\mathfrak{s}^{2}}{p}\left\{\exp\left(\frac{9\rho}{2^{|\bar{m}-\bar{\ell}|/2+2}\mathfrak{s}}\right)-1\right\}\right)\right].$
Note that $\rho\leq\underline{c}_{1}\mathfrak{s}\log\bigl(1+p\log\log(n)/\mathfrak{s}^{2}\bigr)$. Similarly as in the dense regime, we first bound
$\mathsf{E}\left[\exp\left(\frac{\mathfrak{s}^{2}}{p}\left\{\exp\left(\frac{9\rho}{2^{|\bar{m}-\bar{\ell}|/2+2}\mathfrak{s}}\right)-1\right\}\right)\mathbb{I}_{\left\{|\bar{m}-\bar{\ell}|\leq\log\log(n)\right\}}\right]$
$\leq$ $\exp\left(\frac{\mathfrak{s}^{2}}{p}\left\{\exp\left(\frac{9\rho}{4\mathfrak{s}}\right)-1\right\}\right)\mathsf{P}\left\{|\bar{m}-\bar{\ell}|\leq\log\log(n)\right\}$
$\leq$ $\exp\left(\log\log(n)\cdot\frac{\mathfrak{s}^{2}}{p\log\log(n)}\left\{\left(1+\frac{p\log\log(n)}{\mathfrak{s}^{2}}\right)^{{9\underline{c}_{1}}/{4}}-1\right\}\right)\mathsf{P}\left\{|\bar{m}-\bar{\ell}|\leq\log\log(n)\right\}$
$\leq$ $\log(n)^{\alpha_{1}}\cdot\frac{2\log\log(n)+1}{(1-2\nu)\log_{2}(n)-2}\;\to\;0,\qquad\text{as }\;n\to\infty,$
where $\alpha_{1}=\max_{0<x\leq 1}x\bigl((1+x^{-1})^{9\underline{c}_{1}/4}-1\bigr)=2^{9\underline{c}_{1}/4}-1<0.3$. We next bound
$\mathsf{E}\left[\exp\left(\frac{\mathfrak{s}^{2}}{p}\left\{\exp\left(\frac{9\rho}{2^{|\bar{m}-\bar{\ell}|/2+2}\mathfrak{s}}\right)-1\right\}\right)\mathbb{I}_{\left\{|\bar{m}-\bar{\ell}|>\log\log(n)\right\}}\right]$
$\leq$ $\exp\left(\frac{\mathfrak{s}^{2}}{p}\left\{\exp\left(\frac{9\rho}{4\log(n)^{\log(2)/2}\mathfrak{s}}\right)-1\right\}\right)$
$\leq$ $\exp\left(\log\log(n)\cdot\frac{\mathfrak{s}^{2}}{p\log\log(n)}\left\{\left(1+\frac{p\log\log(n)}{\mathfrak{s}^{2}}\right)^{\frac{9\underline{c}_{1}}{4\log(n)^{\log(2)/2}}}-1\right\}\right)$
$\leq$ $\exp\left(\log\log(n)\cdot\left\{2^{\frac{9\underline{c}_{1}}{4\log(n)^{\log(2)/2}}}-1\right\}\right)\;\to\;1,\qquad\text{as }\;n\to\infty,$
where the last inequality follows from the fact that $\max_{0<x\leq 1}x\bigl((1+x^{-1})^{\alpha_{2}}-1\bigr)=2^{\alpha_{2}}-1$ for any $\alpha_{2}\in(0,1)$.
Combining both scenarios, we can have, for sufficiently large $n$,
$\mathsf{E}\left[\left\{1+\frac{\mathfrak{s}}{p}\left(\cosh\left(\frac{9\rho}{2^{|\bar{m}-\bar{\ell}|/2+1}\mathfrak{s}}\right)^{1/2}-1\right)\right\}^{\mathfrak{s}}\right]\leq\frac{4}{3},$
which implies $\chi^{2}(\mathsf{P}^{(1)},\mathsf{P}^{(0)})+1\leq\frac{3}{2}\cdot\frac{4}{3}$, that is, $\chi^{2}(\mathsf{P}^{(1)},\mathsf{P}^{(0)})\leq 1$. Thus
$\frac{\exp\left(-\chi^{2}(\mathsf{P}^{(1)},\mathsf{P}^{(0)})\right)}{4}\geq\frac{1}{4e},$
which is the intended lower bound.
@@ B.1.2 Supporting lemmas
For $\bm{u}\in\mathbb{R}^{p}$, let $\mathsf{P}_{\bm{u}}$ denote the centred $(p+1)$-dimensional Gaussian distribution with covariance matrix
$\begin{bmatrix}\bm{\Sigma}&\bm{u}\\ \bm{u}^{\top}&\bm{u}^{\top}\bm{u}+\sigma^{2}\end{bmatrix}\in\mathbb{R}^{(p+1)\times(p+1)}$
for some positive definite matrix $\bm{\Sigma}$. If $\lvert\bm{\Sigma}^{1/2}\bm{u}\rvert_{2}\cdot\lvert\bm{\Sigma}^{1/2}\bm{v}\rvert_{2}+\langle\bm{\Sigma}^{1/2}\bm{u},\bm{\Sigma}^{1/2}\bm{v}\rangle<\sigma^{2}$, then
$\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\frac{\mathrm{d}\mathsf{P}_{\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\frac{\mathrm{d}\mathsf{P}_{\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)=\left(\left(1-\frac{\langle\bm{\Sigma}^{1/2}\bm{u},\bm{\Sigma}^{1/2}\bm{v}\rangle}{\sigma^{2}}\right)^{2}-\left(\frac{\lvert\bm{\Sigma}^{1/2}\bm{u}\rvert_{2}\cdot\lvert\bm{\Sigma}^{1/2}\bm{v}\rvert_{2}}{\sigma^{2}}\right)^{2}\right)^{-1/2}.$
@@
It follows from a general result for centred multivariate Gaussians (see e.g. Lemma 11 in Cai and Guo, 2017). For completeness, we give below a direct computation:
@@
$=\mathrm{det}\left(\mathbf{I}-\frac{1}{\sigma^{2}}\left(\tilde{\bm{u}}\tilde{\bm{v}}^{\top}+\tilde{\bm{v}}\tilde{\bm{u}}^{\top}\right)\right)^{-1/2}$
$=\left(\left(1-\frac{\langle\tilde{\bm{u}},\tilde{\bm{v}}\rangle}{\sigma^{2}}\right)^{2}-\left(\frac{|\tilde{\bm{u}}|_{2}\cdot|\tilde{\bm{v}}|_{2}}{\sigma^{2}}\right)^{2}\right)^{-1/2},$
where $\tilde{\bm{u}}:=\bm{\Sigma}^{1/2}\bm{u}$, $\tilde{\bm{v}}:=\bm{\Sigma}^{1/2}\bm{v}$ and the last equality can be seen from the fact that
[…]
@@
Note that in Lemma B.2 the condition $|\bm{\Sigma}^{1/2}\bm{u}|_{2}\cdot|\bm{\Sigma}^{1/2}\bm{v}|_{2}+\langle\bm{\Sigma}^{1/2}\bm{u},\bm{\Sigma}^{1/2}\bm{v}\rangle<\sigma^{2}$ is needed to ensure the invertibility of the matrix $\mathbf{I}-\frac{1}{\sigma^{2}}\bm{\Sigma}^{1/2}\left(\bm{u}\bm{v}^{\top}+\bm{v}\bm{u}^{\top}\right)\bm{\Sigma}^{1/2}$. In the special case of $|\bm{\Sigma}^{1/2}\bm{u}|_{2}\cdot|\bm{\Sigma}^{1/2}\bm{v}|_{2}+\langle\bm{\Sigma}^{1/2}\bm{u},\bm{\Sigma}^{1/2}\bm{v}\rangle=\sigma^{2}$, by the same calculation, we have
$\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\frac{\mathrm{d}\mathsf{P}_{\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\frac{\mathrm{d}\mathsf{P}_{\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)=\left(1-\frac{\langle\bm{\Sigma}^{1/2}\bm{u},\bm{\Sigma}^{1/2}\bm{v}\rangle}{\sigma^{2}}+\frac{|\bm{\Sigma}^{1/2}\bm{u}|_{2}\cdot|\bm{\Sigma}^{1/2}\bm{v}|_{2}}{\sigma^{2}}\right)^{-1/2}.$
However, if $|\bm{\Sigma}^{1/2}\bm{u}|_{2}\cdot|\bm{\Sigma}^{1/2}\bm{v}|_{2}+\langle\bm{\Sigma}^{1/2}\bm{u},\bm{\Sigma}^{1/2}\bm{v}\rangle>\sigma^{2}$, we have $\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\tfrac{\mathrm{d}\mathsf{P}_{\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\tfrac{\mathrm{d}\mathsf{P}_{\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)=\infty$.
@@
See either the first proof in Hoeffding (1963, Theorem 4 on pages 21–23), or a more explicit one in Aldous (1985, Proposition (20.6) on pages 171–172), or a systematic treatment in Kemperman (1973, Section 3). ∎